OpenStudy (anonymous):

how do i write csc 2x in terms of csc x? Also sec 2x in form of sec x? Assuming that 0

OpenStudy (anonymous):

I know a way of doing this, but it isn't simple.

OpenStudy (dumbcow):

\[\csc (2x) = \frac{1}{\sin (2x)} = \frac{1}{2 \sin x \cos x}\] \[= \frac{1}{2 \sin x \sqrt{1-\sin^{2} x}}\] \[=\large \frac{\csc x}{2\sqrt{1-\frac{1}{\csc^{2} x}}}\] \[=\frac{\csc^{2} x}{2 \sqrt{\csc^{2} x -1}}\]